# A Novel Garment Transfer Method Supervised by Distilled Knowledge of Virtual Try-on Model

### v1

@@ Abstract
When a shopper chooses garments online, garment transfer technology wears the garment from the model image onto the shopper’s image, allowing the shopper to decide whether the garment is suitable for them. As garment transfer leverages wild and cheap person image as garment condition, it has attracted tremendous community attention and holds vast commercial potential. However, since the ground truth of garment transfer is almost unavailable in reality, previous studies have treated garment transfer as either pose transfer or garment-pose disentanglement, and trained garment transfer in self-supervised learning, yet do not cover garment transfer intentions completely. Therefore, the training supervising the garment transfer is a rock-hard issue. Notably, virtual try-on technology has exhibited superior performance using self-supervised learning. We supervise the garment transfer training via knowledge distillation from virtual try-on. Specifically, we first train the transfer parsing reasoning model at multi-phases to provide shape guidance for downstream tasks. The transfer parsing reasoning model learns the response and feature knowledge from the try-on parsing reasoning model and absorbs the hard knowledge from the ground truth. By leveraging the warping knowledge from virtual try-on, we estimate a progressive flow to precisely warp the garment by learning the shape and content correspondence. To enhance transfer realism, we propose a well-designed arm regrowth task to infer exposed skin pixel content. Experiments demonstrate that our method has state-of-the-art performance in transferring garments between person compared with other virtual try-on and garment transfer methods.
@@ 1 Introduction
Virtual try-on wears the garment-in-shop into a person, while garment transfer wears the garment from one person to another. As Fig. 1 shows, the garment-in-shop of virtual try-on is captured in a controlled environment using professional techniques, while the person image of garment transfer is wilder and cheaper. In a word, garment transfer also allows shoppers to perceive the garment-wearing effect but does not require specific and expensive garment conditions. However, the pair of garment-in-shop and the try-on result is easy to collect, and the pair of the person image and the transfer result is almost unavailable. Thus, garment transfer is a user-friendly but training-hard topic. As virtual try-on is implemented by self-supervised learning and both technologies share the same purpose, we propose to supervise the training of garment transfer by distilling knowledge from virtual try-on. To the best of our knowledge, this is the first time to implement garment transfer via knowledge distillation. In this paper, we present a novel garment transfer pipeline and knowledge distillation process by exploiting previous virtual try-on studies.
@@
To wear a garment-in-shop, the virtual try-on method warps the garment-in-shop towards the person’s pose and preserves the other person’s features. It involves many topics, such as image warping, image synthesis, and image fusion. VITON (Han et al., 2018) proposed the concept of image-based virtual try-on for the first time and provided an inspiring pipeline to implement it. It synthesizes a coarse try-on result, warps the garment-in-shop toward this result, and combines the synthesized and warped results to refine the garment texture. Alternatively, CP-VTON (Wang et al., 2018) offers a different pipeline where the warping and synthesis sequences are reversed. Since the paired garment-in-shop and the try-on result are available, they were both trained in a self-supervised manner by wearing a garment-in-shop onto a person dressed in the same garment. This self-supervision learning enables models to develop warping and rendering capacities that can handle new garment-wearing cases during the testing phase.
@@
Intuitively, the garment input of virtual try-on is nearly aligned and has a high pixel ratio since it is captured at a hanging status in the shop. Conversely, the garment in the person image is entangled with the pose feature of the person, presenting challenges for the garment transfer model. When two persons wear different garments in various poses, the ideal ground truth for garment transfer is for one person to wear the other person’s garment and maintain its intrinsic pose. However, it is almost impossible to collect such ground truth data in reality. Therefore, the primary focus of garment transfer studies is on how to supervise its learning, and there are two dominant ways to achieve this. One approach (Raj et al., 2018) is to disentangle the garment in person image from pose and train the model in a self-supervised manner to reconstruct the person image conditioned on the spatial-agnostic garment and pose features. The other approach (Yang and Lin, 2021) is to directly alter the pose feature of the garment by warping. With a pair of person images where a person wears the same garment but has different poses, the model is trained to warp the garment from one pose to another.
Since pair fashion data is essential for the training of virtual try-on and garment transfer, several public datasets have been released for research purposes. The Zalando dataset (Han et al., 2018) contains the paired garment-in-shop and try-on result images in a resolution of 256×256. A similar virtual try-on dataset (Choi et al., 2021) is collected but in a higher resolution of 1024×768. Deepfashion2 (Ge et al., 2019) provided paired person images with different poses in the in-shop clothes retrieval benchmark, which has been extensively employed for self-supervised learning of garment transfer and pose transfer issues.
@@
With the easy-to-access dataset and easy-to-follow pipeline, virtual try-on studies are more flourishing than garment transfer studies. Existing garment transfer studies encounter several issues that hinder their ability to achieve transfer results comparable to virtual try-on studies. 1) The self-supervised learning does not completely conform to its intention. The disentanglement of garment and pose is not imperative to garment transfer and impose a burden on models in reconstructing high-frequency details when re-entangling the garment and pose. These studies typically train garment transfer as pose transfer within the same person, leading to shape and style preservation issues when transferring garments to other person; 2) The task execution sequence is unreasonable. As the traditional learning methods cannot supervise the training of transfer parsing reasoning conditioned on transferred garment and person image, transfer parsing reasoning only follows garment warping and guides final image synthesis. Without transfer parsing prior, garment transfer guided by target pose is inevitable to yield misalignment and artifacts. 3) A specialized preservation mechanism for person features is missing. Both direct propagation and content inference are required in transferring the person’s features, such as the face and skin texture, into the final result. However, they only consider the direct composition of the warped garment and person image without designing specialized mechanisms to handle the multi-tasks involved in person feature preservation.
@@
(1) Teaching transfer parsing reasoning in multi-phase. The transfer parsing reasoning model learns the response and feature knowledge from the pre-trained try-on parsing reasoning model in garment transfer. We then transfer the garment back to its originator so that the transfer parsing reasoning model absorbs the hard knowledge from the ground truth to improve its robustness.
@@
(3) Experiments demonstrate that our method has state-of-the-art performance compared other the virtual try-on and garment transfer methods in garment transfer. Our method performs well in transfer parsing reasoning, garment warping, new skin inference, and final result synthesis whatever the dressing and pose of person images are.
@@ 2 Related Work
Virtual Try-on VITON (Han et al., 2018) and CP-VTON (Wang et al., 2018) determined the fundamental strategy to implement virtual try-on, and the subsequent studies focused on optimizing garment warping and person reservation. Yang H et al. introduced second-order constraint into the TPS deformation in ACGPN (Yang et al., 2020). Minar M R et al. proposed CP-VTON+ (Minar et al., 2020) extended on CP-VITON where the skin label is updated and the garment-in-shop mask is concatenated into the image synthesis. Neuberger A et al. (Neuberger et al., 2020) proposed an online optimization step to optimize fine details reservation. Ge C et al. introduced cycle consistency to implement self-supervision learning in DCTON (Ge et al., 2021a). However, the coarse-grained TPS deformation used by these methods inevitably results in misalignment between the garment and person, requiring warping and rendering to be combined, which can introduce artifacts into the final result. To address this issue, some studies have used optical flow with higher DOF for finer warping. Ge Y et al. proposed PF-AFN (Ge et al., 2021b) to estimate appearance flow from coarse to fine. He S et al. (He et al., 2022) proposed Flow Style VTON (FSV) to estimate a global warping tendency via style modulation based on PF-AFN. Additionally, some studies have used knowledge distillation to make their models independent of parsing and pose. Issenhuth T et al. proposed WUTON (Issenhuth et al., 2020) to transfer warping knowledge and disambiguate the human parser error influence. PF-AFN (Ge et al., 2021b) and FSV (He et al., 2022) linked the teacher model output with the student model input to implement the parser-free virtual try-on.
@@
Garment Transfer It is regarded as garment-pose disentanglement or pose transfer for self-supervised learning. For the former, as the originator of garment transfer, SwapNet (Raj et al., 2018) learned the categories embedding by ROI pooling to disentangle the garment with the pose. Xie Z et al. proposed PASTA-GAN (Xie et al., 2021) to cut the garment into multiple patches for the disentanglement and exploited two StyleGAN2 to re-entangle garment patches with pose while considering style and texture. For the latter, Yang F et al. proposed two complementary warping in CT-Net (Yang and Lin, 2021) to implement garment pose changing. Liu T et al. proposed SPATT (Liu et al., 2021) to establish a spatial correspondence in UV space to facilitate garment warping and unobserved region inference.
@@
Knowledge Distillation It was proposed by Hinton G et al. in (Hinton et al., 2015), where the pre-trained teacher model supervises the student model learning by teaching various knowledge. This approach aims to compress the parameter quantity and training data, enabling the student model to be lightweight. The distilled knowledge can be categorized into three types: response knowledge, feature knowledge, and relation knowledge. Specifically, response knowledge (Hinton et al., 2015) and feature knowledge (Zagoruyko and Komodakis, 2017) are extracted at the last and hint layer, containing prediction and processing information. And relation knowledge (Ding et al., 2022) describes the relationship between data or feature maps. The methods of knowledge distillation are categorized into offline distillation, online distillation, and self-distillation. Offline distillation employs a large model pre-trained on a large dataset as the teacher model and teaches a lightweight student model. In contrast, online distillation (Liu et al., 2022) synchronously trains the teacher and student models to achieve better knowledge transfer when the large model is unavailable for specific tasks. Because of its universality and flexibility, knowledge distillation has been applied in several domains, including multi-view learning (Tian et al., 2022) and speaker extraction (Huang et al., 2022).
@@ 3.1 Outline
As Fig. 2 shows, we exploit the same pipeline for virtual try-on and garment transfer: parsing reasoning, garment warping, arm regrowth, and remaining content propagation. Virtual try-on teaches garment transfer via knowledge distillation in the parsing reasoning and garment warping tasks. For virtual try-on, the teacher transfer parsing reasoning model ${{{\cal T}_{p}}}$ and teacher flow warping model ${{{\cal T}_{w}}}$ are trained in a self-supervised manner. In the testing phase, ${{{\cal T}_{p}}}$ reasons the try-on parsing ${{{\cal M}_{TAB}}}$ where person ${A}$ wears the garment-in-shop ${B}$. Guided by the garment mask ${{\cal M}_{TAB}^{hc}}$ of ${{{\cal M}_{TAB}}}$, ${{{\cal T}_{w}}}$ warps ${{\cal I}_{B}^{c}}$ towards ${{\cal M}_{TAB}^{hc}}$ to yield ~${\tilde{\cal I}_{TAB}^{hc}}$. For garment transfer, the student transfer parsing reasoning model ${{{\cal S}_{p}}}$ and the student flow warping model ${{{\cal S}_{w}}}$ are trained with supervision information distilled from ${{{\cal T}_{p}}}$ and ${{{\cal T}_{w}}}$. ${{{\cal S}_{p}}}$ reasons the transfer parsing ${{{\cal M}_{SAB}}}$ where person ${A}$ wears the garment of person ${B}$ ( Sec. 3.2). With this target shape ${{\cal M}_{SAB}^{hc}}$, ${{{\cal S}_{w}}}$ warps garment ${{\cal I}_{B}^{hc}}$ to yield ~${\tilde{\cal I}_{SAB}^{hc}}$ ( Sec. 3.3). For the transfer realism, the arm regrowth model ${{G_{r}}}$ reasons more upper body skin ~${\tilde{\cal I}_{SAB}^{ha}}$ when ${{\cal M}_{SAB}^{ha}>{\cal M}_{A}^{ha}}$ ( Sec. 3.4). And the remaining content of person ${A}$ is directly propagated by multiplying ${{\cal M}_{SAB}^{hr}}$ with ${{{\cal I}_{A}}}$. In the end, we concatenate ~${\tilde{\cal I}_{SAB}^{ha}}$, ~${\tilde{\cal I}_{SAB}^{hr}}$, and ~${\tilde{\cal I}_{SAB}^{hc}}$ to obtain ~${\tilde{\cal I}_{SAB}}$.
@@ 3.2.1 Parsing Reasoning Insight
Garment transfer is a complex computer vision task that attempts to transfer garments from person ${B}$ to another person ${A}$, involving various shape-level and content-level tasks. To alleviate the computational burden of a single model in performing these multi-level tasks, we prioritize the transfer parsing reasoning at the shape level, as it offers the critical shape guidance required for downstream tasks. This transfer parsing has the garment-agnostic features of person ${A}$ and the transferred garment features of person ${B}$.
The fundamental challenge of transfer parsing reasoning lies in how to effectively entangle the garment feature of person ${B}$ with the pose feature of person ${A}$, such that the garment conforms to the global pose feature of person ${A}$ while preserving its original style and category. This presents a significant challenge for both traditional fully supervised learning and self-supervised learning approaches. In the case of fully supervised learning, acquiring ground truth data where person ${A}$ wears the garment of person ${B}$ is essential, but such data is rarely available in practice. Alternatively, in the case of self-supervised learning, the garment features undergo pose augmentation and the model is trained to entangle this augmented garment feature with its original pose feature. However, as data augmentation alone is insufficient to disentangle the original pose feature from the garment, the entanglement learning is prone to degenerate the plain mapping, leading to unreliable reasoning during the testing phase.
@@ 3.2.2 Teacher Try-on Parsing Reasoning Model
We exploit the pre-trained Grapy-ML (He et al., 2020) to extract person parsing ${{{\cal M}_{A}}}$, ${{{\cal M}_{B}}}$ from images ${{{\cal I}_{A}}}$, ${{{\cal I}_{B}}}$. To facilitate subsequent processing, the person parsing categories are grouped into background, hair, face, upper body skin, upper garment, lower garment, and thigh. Additionally, the pre-trained Densepose model (Güler et al., 2018) is utilized to extract the dense pose features ${{{\cal P}_{A}}}$, ${{{\cal P}_{B}}}$ from images ${{{\cal I}_{A}}}$, ${{{\cal I}_{B}}}$. We adopt U2-Net as the backbone of ${{{\cal T}_{p}}}$ , which is proficient in capturing multiscale contextual information and is well-suited for the given reasoning task.
As described in Sec. 3.2.1, the objective of ${{{\cal T}_{p}}}$ is to entangle shape features of a new garment-in-shop with pose features of the person while also propagating garment-agnostic features of the person. Achieving this necessitates providing independent priors for each feature. Accordingly, we leverage ${{\cal M}_{A}^{c}}$ and ${{{\cal P}_{A}}}$ to represent the garment shape and person pose, respectively. Since the upper body skin and background categories entangle with the garment category, including them in garment-agnostic features can hinder entanglement learning. Therefore, we exclude these categories from the parsing ${{{\cal M}_{A}}}$ and merge the remaining ones into ${{\cal M}_{A}^{hr}}$, which depicts the garment-agnostic shape in person parsing. During the self-supervised training, we concatenate ${{\cal M}_{A}^{c}}$,${{{\cal P}_{A}}}$,${{\cal M}_{A}^{hr}}$ and feed them into the model ${{{\cal T}_{p}}}$ to predict ~${{\tilde{\cal M}_{A}}}$, which is supervised with the ground truth ${{{\cal M}_{A}}}$. During testing, we substitute ${{\cal M}_{A}^{c}}$ with ${{\cal M}_{B}^{c}}$ to predict ${{{\cal M}_{TAB}}}$, where person ${A}$ wears the garment-in-shop ${B}$. We denote this process as ${{{\cal M}_{TAB}}={{\cal T}_{p}}\left({{\cal M}_{B}^{c},{{\cal P}_{A}},{\cal M}_{A}^{hr}}\right)}$ in the formula.

### v2

@@ Abstract
Garment transfer can wear the garment of the model image onto the personal image. As garment transfer leverages wild and cheap garment input, it has attracted tremendous attention in the community and has a huge commercial potential. Since the ground truth of garment transfer is almost unavailable in reality, previous studies have treated garment transfer as either pose transfer or garment-pose disentanglement, and trained garment transfer in self-supervised learning, However, these implementation methods do not cover garment transfer intentions completely and face the robustness issue in the testing phase. Notably, virtual try-on technology has exhibited superior performance using self-supervised learning, we propose to supervise the garment transfer training via knowledge distillation from virtual try-on. Specifically, the overall pipeline is first to infer a garment transfer parsing, and to use it to guide downstream warping and inpainting tasks. The transfer parsing reasoning model learns the response and feature knowledge from the try-on parsing reasoning model and absorbs the hard knowledge from the ground truth. The progressive flow warping model learns the content knowledge from virtual try-on for a reasonable and precise garment warping. To enhance transfer realism, we propose an arm regrowth task to infer exposed skin. Experiments demonstrate that our method has state-of-the-art performance in transferring garments between persons compared with other virtual try-on and garment transfer methods.
@@ 1 Introduction
Virtual try-on wears the garment-in-shop into a person, while garment transfer wears the garment from one person to another. As Fig. 1 shows, the garment-in-shop of virtual try-on is captured in a controlled environment with professional equipment, while the personal image of garment transfer is wilder and cheaper. In a word, garment transfer also allows shoppers to perceive the garment-wearing effect but does not require specific and expensive garment conditions. However, the pair of garment-in-shop and the try-on result is easy to collect, and the pair of the person image and the transfer result is almost unavailable. Thus, garment transfer is a user-friendly but training-hard topic. As virtual try-on can be implemented by self-supervised learning and both technologies share the same purpose, we propose to supervise the training of garment transfer by distilling knowledge from virtual try-on. To the best of our knowledge, this is the first time to implement garment transfer via knowledge distillation. In this paper, we present a novel garment transfer pipeline and knowledge distillation process by exploiting previous virtual try-on studies.
@@
To wear a garment-in-shop, the virtual try-on method warps the garment-in-shop towards the personal pose and preserves the other personal features. It involves many topics, such as image deformation, image inpainting, and image fusion. VITON [1] proposed the concept of image-based virtual try-on for the first time and provided an inspiring implementation pipeline. It synthesizes a coarse try-on result, warps the garment-in-shop toward this result, and combines the synthesized and warped results to refine the garment texture. Alternatively, CP-VTON [2] offers a different pipeline where the warping and synthesis sequences are reversed. Since the paired garment-in-shop and the try-on result are available, they were both trained in a self-supervised manner by wearing a garment-in-shop onto a person dressed in the same garment. This self-supervision learning enables models to have perfect warping and inpainting capacities for general cases.
@@
Intuitively, the garment-in-shop is nearly aligned and has a high pixel ratio since it is captured at a hanging status. Conversely, the garment in the person image is entangled with the personal pose, presenting challenges for the garment transfer model. When two persons wear different garments with different poses, the ground truth of garment transfer is one person wears the garment of another person while maintaining itself pose. However, it is almost impossible to collect such ground truth in reality. Therefore, the research key point of garment transfer is how to supervise its learning, and there are two dominant ways. One approach [3] is to disentangle the garment and pose in the personal image and re-entangle the pose and spatial-agnostic garment to generate the personal image. The other approach [4] is to directly alter the pose feature in the garment by warping. With a pair of person images where a person wears the same garment but has different poses, the model is trained to warp the garment from one pose to another.
@@
With the easy-to-access dataset and easy-to-follow pipeline, virtual try-on studies are more flourishing than garment transfer studies. Existing garment transfer studies encounter several issues that hinder their ability to achieve transfer results comparable to virtual try-on studies. 1) The implementation method does not completely conform to its intention. It is imperative to disentangle pose from garment and results in high-frequency details in re-entangling the garment and pose. Training garment transfer as pose transfer within the same person faces shape and style preservation issues in transferring a garment to a different person; 2) The multi-task execution sequence is unreasonable. Garment warping is guided by a target pose instead of a transfer parsing, while the transfer parsing is reasoned after the garment warping task to guide final image synthesis. It is inevitable to yield misalignment and artifacts. 3) The specialized preservation mechanism of personal features receives inadequate study. It needs direct propagation and content inference to preserve the face and upper body skin. They just combine the warped garment and source personal features and result in artifacts in the new exposure part without a specialized inference mechanism.
@@
(1) Teaching transfer parsing reasoning in multi-phase. The transfer parsing reasoning model learns the response and feature knowledge from the pre-trained try-on parsing reasoning model. We then transfer the garment back to its originator so that the transfer parsing reasoning model absorbs the hard knowledge from the ground truth to improve its robustness.
@@
(3) Experiments demonstrate that our method has state-of-the-art performance compared to other virtual try-on and garment transfer methods in garment transfer. Our method performs well in transfer parsing reasoning, garment warping, new skin inference, and final result synthesis whatever the dressing and pose of person images are.
@@ 2 Related Work
Virtual Try-on VITON [1] and CP-VTON [2] determined the fundamental strategy to implement virtual try-on, and the subsequent studies focused on optimizing garment warping and person reservation. Yang H et al. introduced second-order constraint into the TPS deformation in ACGPN [5]. Minar M R et al. proposed CP-VTON+ [6] extended on CP-VITON where the skin label is updated and the garment-in-shop mask is concatenated into the image synthesis. Neuberger A et al. [7] proposed an online optimization step to optimize fine details reservation. Ge C et al. introduced cycle consistency to implement self-supervision learning in DCTON [8]. However, the coarse-grained TPS deformation used by these methods inevitably results in misalignment between the garment and person, requiring warping and rendering to be combined, which can introduce artifacts into the final result. To address this issue, some studies have used optical flow with higher DOF for finer warping. Ge Y et al. proposed PF-AFN [9] to estimate appearance flow from coarse to fine. Fang N et al. [10] proposed a warping-mapping-composition mechanism. He S et al. [11] proposed Flow Style VTON (FSV) to estimate a global warping tendency via style modulation based on PF-AFN. Additionally, some studies have used knowledge distillation to make their models independent of parsing and pose. Issenhuth T et al. proposed WUTON [12] to transfer warping knowledge and disambiguate the human parser error influence. PF-AFN [9] and FSV [11] linked the teacher model output with the student model input to implement the parser-free virtual try-on.
@@
Garment Transfer It is regarded as garment-pose disentanglement or pose transfer for self-supervised learning. For the former, as the originator of garment transfer, SwapNet [3] learned the categories embedding by ROI pooling to disentangle the garment with the pose. Xie Z et al. proposed PASTA-GAN [13] to cut the garment into multiple patches for the disentanglement and exploited two StyleGAN2 to re-entangle garment patches with pose while considering style and texture. For the latter, Yang F et al. proposed two complementary warping in CT-Net [4] to implement garment pose changing. Liu T et al. proposed SPATT [14] to establish a spatial correspondence in UV space to facilitate garment warping and unobserved region inference.
@@
Knowledge Distillation It was proposed by Hinton G et al. in [15], where the pre-trained teacher model supervises the student model learning by teaching various knowledge. This approach aims to compress the parameter quantity and training data, enabling the student model to be lightweight. The distilled knowledge can be categorized into three types: response knowledge, feature knowledge, and relation knowledge. Specifically, response knowledge [15] and feature knowledge [16] are extracted at the last and hint layer, containing prediction and processing information. And relation knowledge [17] describes the relationship between data or feature maps. The methods of knowledge distillation are categorized into offline distillation, online distillation, and self-distillation. Offline distillation employs a large model pre-trained on a large dataset as the teacher model and teaches a lightweight student model. In contrast, online distillation [18] synchronously trains the teacher and student models to achieve better knowledge transfer when the large model is unavailable for specific tasks. Because of its universality and flexibility, knowledge distillation has been applied in several domains, including multi-view learning [19] and speaker extraction [20].
@@ 3.1 Outline
As Fig. 2 shows, we exploit the same pipeline for virtual try-on and garment transfer: parsing reasoning, garment warping, arm regrowth, and remaining content propagation. Virtual try-on teaches garment transfer via knowledge distillation in the parsing reasoning and garment warping tasks. For virtual try-on, the teacher transfer parsing reasoning model ${{{\cal T}_{p}}}$ and teacher flow warping model ${{{\cal T}_{w}}}$ are trained in a self-supervised manner. In the testing phase, ${{{\cal T}_{p}}}$ reasons the try-on parsing ${{{\cal M}_{TAB}}}$ where person ${A}$ wears the garment-in-shop ${B}$. Guided by the garment mask ${{\cal M}_{TAB}^{hc}}$ of ${{{\cal M}_{TAB}}}$, ${{{\cal T}_{w}}}$ warps ${{\cal I}_{B}^{c}}$ towards ${{\cal M}_{TAB}^{hc}}$ to yield ~${\tilde{\cal I}_{TAB}^{hc}}$. For garment transfer, the student transfer parsing reasoning model ${{{\cal S}_{p}}}$ and the student flow warping model ${{{\cal S}_{w}}}$ are trained with supervision information distilled from ${{{\cal T}_{p}}}$ and ${{{\cal T}_{w}}}$. ${{{\cal S}_{p}}}$ reasons the transfer parsing ${{{\cal M}_{SAB}}}$ where person ${A}$ wears the garment of person ${B}$ ( Sec. 3.2). With this target shape ${{\cal M}_{SAB}^{hc}}$, ${{{\cal S}_{w}}}$ warps garment ${{\cal I}_{B}^{hc}}$ to yield ~${\tilde{\cal I}_{SAB}^{hc}}$ ( Sec. 3.3). For the transfer realism, the arm regrowth model ${{G_{r}}}$ reasons more upper body skin ~${\tilde{\cal I}_{SAB}^{ha}}$ when ${{\cal M}_{SAB}^{ha}>{\cal M}_{A}^{ha}}$ ( Sec. 3.4). And for the remaining content, we just need inherit its pixel-level content from person ${A}$, via Boolean multiplying ${{\cal M}_{SAB}^{hr}}$ with ${{{\cal I}_{A}}}$. In the end, we concatenate ~${\tilde{\cal I}_{SAB}^{ha}}$, ~${\tilde{\cal I}_{SAB}^{hr}}$, and ~${\tilde{\cal I}_{SAB}^{hc}}$ to obtain ~${\tilde{\cal I}_{SAB}}$.
@@ 3.2.1 Parsing Reasoning Insight
Garment transfer is a complex computer vision task that attempts to transfer garments from person ${B}$ to another person ${A}$, involving various shape-level and content-level tasks. To alleviate the computational burden of a single model to conduct these multi-level tasks, we prioritize the transfer parsing reasoning at the shape level, as it offers the critical shape guidance required for downstream tasks. This transfer parsing has the garment-agnostic features of person ${A}$ and the transferred garment features of person ${B}$.
The fundamental challenge of transfer parsing reasoning lies in how to effectively entangle the garment feature of person ${B}$ with the pose feature of person ${A}$, such that the garment conforms to the global pose feature of person ${A}$ while preserving its original style and category. This presents a significant challenge for both traditional fully supervised learning and self-supervised learning approaches. In the case of fully supervised learning, it is essential to acquire ground truth where person ${A}$ wears the garment of person ${B}$, but such data is rarely available in reality. Alternatively, in the case of self-supervised learning, the garment features undergo pose augmentation and the model is trained to entangle this augmented garment feature with its original pose feature. However, as data augmentationis insufficient to disentangle the original pose feature from the garment, the entanglement learning is prone to degenerate the plain mapping, leading to unreliable reasoning during the testing phase.
@@ 3.2.2 Teacher Try-on Parsing Reasoning Model
We exploit the pre-trained Grapy-ML [21] to extract person parsing ${{{\cal M}_{A}}}$, ${{{\cal M}_{B}}}$ from images ${{{\cal I}_{A}}}$, ${{{\cal I}_{B}}}$. To facilitate subsequent processing, the person parsing categories are grouped into background, hair, face, upper body skin, upper garment, lower garment, and thigh. Additionally, the pre-trained Densepose model [22] is utilized to extract the dense pose features ${{{\cal P}_{A}}}$, ${{{\cal P}_{B}}}$ from images ${{{\cal I}_{A}}}$, ${{{\cal I}_{B}}}$. We adopt U2-Net as the backbone of ${{{\cal T}_{p}}}$ , which is proficient in capturing multiscale contextual information and is well-suited for the given reasoning task.
As described in Sec. 3.2.1, the objective of ${{{\cal T}_{p}}}$ is to entangle shape features of a new garment-in-shop with pose features of the person while also propagating garment-agnostic features of the person. It necessitates providing independent priors for each feature for this task. Accordingly, we leverage ${{\cal M}_{A}^{c}}$ and ${{{\cal P}_{A}}}$ to represent the garment shape and person pose, respectively. Since the upper body skin and background categories entangle with the garment category, including them in garment-agnostic features will hinder entanglement learning. Therefore, we exclude these categories from the parsing ${{{\cal M}_{A}}}$ and merge the remaining ones into ${{\cal M}_{A}^{hr}}$, which depicts the garment-agnostic shape in person parsing. During the self-supervised training, we concatenate ${{\cal M}_{A}^{c}}$,${{{\cal P}_{A}}}$,${{\cal M}_{A}^{hr}}$ and feed them into the model ${{{\cal T}_{p}}}$ to predict ~${{\tilde{\cal M}_{A}}}$, which is supervised with the ground truth ${{{\cal M}_{A}}}$. During testing, we substitute ${{\cal M}_{A}^{c}}$ with ${{\cal M}_{B}^{c}}$ to predict ${{{\cal M}_{TAB}}}$, where person ${A}$ wears the garment-in-shop ${B}$. We denote this process as ${{{\cal M}_{TAB}}={{\cal T}_{p}}\left({{\cal M}_{B}^{c},{{\cal P}_{A}},{\cal M}_{A}^{hr}}\right)}$ in the formula.